OpenStudy (anonymous):

What is the sum of the first 15 terms of the equation an=10n+21. The answer is NOT 1374.

OpenStudy (jdoe0001):

\(\bf S_{\color{red}{ n}}=\cfrac{{\color{red}{ n}}(a_1+a_{\color{red}{ n}})}{2}\qquad \qquad S_{\color{red}{ 15}}=\cfrac{{\color{red}{ 15}}(a_1+a_{\color{red}{ 15}})}{2}\)

OpenStudy (anonymous):

Would the beginning of the sequence look like 1,31 ... or 31, 41 ...?

OpenStudy (anonymous):

Like what is a1 supposed to be?

OpenStudy (anonymous):

just put n=1 in the An formula, u get a1=10*1+21=31 then put n=15 in the An forumla, u get a15=10*15+21 then u use Sn=(a1+a15)*15*0.5

OpenStudy (anonymous):

= 1515

OpenStudy (jdoe0001):

if you want to know the beggining of the formula or I gather you mean the 1st term of the sequence just set n = 1

OpenStudy (jdoe0001):

\(\bf a_n=10n+21\) is an arithmetic sequence, it's just already combined including the common difference and "n" and 1st term, in a simplified version

OpenStudy (anonymous):

ok thanks guys!